# N. Deng & Y. Wang & W. Cao

### v1

@@ 5.1 Existence and uniqueness of invariant measure for the numerical solution
where $\nu(\cdot):=\lambda\left(\cdot\cap B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\cap D_{\alpha}\right)/\lambda\left(B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\cap D_{\alpha}\right).$ Letting $\eta=\gamma_{1}\gamma_{2}\lambda\left(B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\cap D_{\alpha}\right)$, we obtain $P_{k^{*}}(\vec{v},B\cap D_{\alpha})\geq\eta\nu(B\cap D_{\alpha})$ for all $B\in\mathscr{B}\big(\mathbb{R}^{N}\big)$ and $\vec{v}\in S\cap D_{\alpha}$. Since $B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\subset S$, it follows that $\nu(S\cap D_{\alpha})=1$ and $\nu\left(\left(S\cap D_{\alpha}\right)^{c}\right)=0$. This completes the proof.
Next we introduce the Lyapunov condition. {assumption} (Lyapunov condition) There is a function $V:\mathbb{R}^{N}\rightarrow[1,\infty)$ with $\lim_{|\vec{v}|\rightarrow\infty}V(\vec{v})=\infty$ and real numbers $\alpha_{1}\in$ $(0,1),\alpha_{2}\in[0,\infty)$ such that
@@
(Mattingly, Stuart & Higham, 2002, Lemma A.6) Suppose $\{\vec{v}_{k}\}_{k\in\mathbb{N}}$ and $\{\vec{w}_{k}\}_{k\in\mathbb{N}}$ are two Markov chains satisfying Assumption 5.2. Let $\gamma\in(\sqrt{\alpha_{1}},1)$ and define
@@
By Lemmas 5.1 and 5.3, we can establish the following theorem.
@@
$\sup_{k\geq 0}\mathbb{E}\left[V\left(\vec{v}_{k}\right)\right]<\infty.$ (80)
@@
It follows from (80) that the sequence $\left\{\pi^{\tau,N}_{m}\right\}_{m\in\mathbb{N}}$ is tight. Consequently, by the Krylov–Bogoliubov Theorem (Hong & Wang, 2019, Theorem 1.2), the Markov chain $\left\{\vec{v}_{k}\right\}_{k\in\mathbb{N}}$ admits an invariant measure $\pi^{\tau,N}$ .
@@
Algorithm 1 Construction of the new Markov chain $\left\{\vec{w}_{k}\right\}_{k\in\mathbb{N}}$
@@
2:Output: Markov chain $\{\vec{w}_{k}\}_{k\in\mathbb{N}}$
@@
4:  Sample $\hat{\psi}_{k}\in\{0,1\}$ with $\mathbb{P}(\hat{\psi}_{k}=1)=1-\eta$ and $\mathbb{P}(\hat{\psi}_{k}=0)=\eta$
5:  Sample $\vec{w}_{k}^{*}\sim\tilde{P}_{k^{*}}(\vec{w}_{k},\cdot)$ and $\vec{\xi}_{k}\sim\nu(\cdot)$ independently
6:  if $\vec{w}_{k}\in S$ then
7:   $\vec{w}_{k+1}\leftarrow\hat{\psi}_{k}\,\vec{w}_{k}^{*}+(1-\hat{\psi}_{k})\,\vec{\xi}_{k}$
8:  else
9:   $\vec{w}_{k+1}\leftarrow\vec{w}_{k}^{*}$
10:  end if
11:end for
@@
This shows that $\pi^{\tau,N}=\hat{\pi}^{\tau,N}$, and hence uniqueness is proved.
@@
Suppose that Assumptions 2-2 hold, then $\left\{u^{\tau,N}(t_{m},\cdot)\right\}_{m\in\mathbb{N}}$ admits a unique invariant measure $\pi^{\tau,N}$.
Proof 5.7.
Since $u^{\tau,N}(t,x)=P_{N}U^{\tau,N}_{t}$, it is sufficient to establish the existence and uniqueness of an invariant measure for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$. From (55), it follows that $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ forms a time-homogeneous Markov chain. According to Theorem 5.4, establishing the existence and uniqueness of an invariant measure for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ necessitates verifying that this Markov chain satisfies both the Lyapunov and minorization conditions restricted to $D_{\alpha}$.
@@
for any $\vec{v}\in l^{2}_{N}$. By applying (55) and using properties of conditional expectation, we obtain
$\mathbb{E}\left[V\left(U^{\tau,N}_{t_{m+1}}\right)\mid\mathscr{F}_{t_{m}}\right]=$ $\sum_{j=1}^{N-1}\lambda_{N,j}^{-1}\mathbb{E}\left[\langle e^{-A_{N}^{2}\tau}U^{\tau,N}_{t_{m}}+\tau A_{N}e^{-A_{N}^{2}\tau}\tilde{F}_{N}(U^{\tau,N}_{t_{m}})+\sigma e^{-A_{N}^{2}\tau}\Delta\beta_{m},\phi_{N,j}\rangle^{2}_{l^{2}_{N}}\mid\mathscr{F}_{t_{m}}\right]$ (81)
@@
Substituting these inequalities into (81), we conclude that
@@
Now, for any $\vec{v}\in S\cap D_{\alpha}$, $\vec{v}^{*}\in\text{Int}(S)\cap D_{\alpha}$, $\delta>0$ and $\vec{w}\in B_{\delta}\left(\vec{v}^{*}\right)\cap D_{\alpha}$, we apply (55) together with the fact that $\left\{\phi_{N,j}\right\}_{j=0}^{N-1}$ is an orthonormal basis of $l^{2}_{N}$ to obtain
@@
This demonstrates that one can choose $\Delta\beta_{m}$ such that $U^{\tau,N}_{t_{m+1}}=\vec{w}$ starting from $U^{\tau,N}_{t_{m}}=\vec{v}$. Since $\left\{\left\langle\Delta\beta_{m},\phi_{N,j}\right\rangle_{l^{2}_{N}}\right\}_{j=1}^{N-1}$ are Gaussian random variables, it follows that $P_{1}\left(\vec{v},B_{\delta}\left(\vec{v}^{*}\right)\cap D_{\alpha}\right)>0$, which verifies the second part of the minorization condition. To verify the third condition in the minorization condition, observe that each Gaussian variable $\left\langle\Delta\beta_{m},\phi_{N,j}\right\rangle_{l^{2}_{N}}$ admits a $C^{\infty}$ density function for $j=1,\cdots,N-1$, and $\left\langle\Delta\beta_{m},\phi_{N,0}\right\rangle_{l^{2}_{N}}=0$. Hence, the one-step transition kernel $P_{1}(\vec{v},\cdot)$ admits a density $p_{1}(\vec{v},\vec{w})$, which is jointly continuous on $(\vec{v},\vec{w})\in\left(S\cap D_{\alpha}\right)\times\left(S\cap D_{\alpha}\right)$. Finally, the time-homogeneity of the Markov chain $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ implies that the density functions $\left\{p_{m}(\vec{v},\vec{w})\right\}_{m\in\mathbb{N}}$ are jointly continuous. This completes the proof.
5.2 Approximation of invariant measure
@@
Suppose that Assumptions 2, 2, 4.2 hold, then for any bounded uniformly continuous function $\phi:\ H\to\mathbb{R}$, there exists a constant $C>0$, independent of $h$ and $\tau$, such that
$\left|\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v)-\int_{V^{N}}\phi(v)\tilde{\pi}^{\tau,N}(\mathrm{d}v)\right|\leq C\left(h^{1-\epsilon}+\tau^{\frac{3}{8}-\frac{\epsilon}{2}}\right),$ (82)
@@
$\lim_{M\rightarrow\infty}\frac{1}{M}\sum_{m=0}^{M-1}\mathbb{E}\left[\phi(u^{\tau,N}(t_{m},\cdot))\right]=\int_{V^{N}}\phi(v)\pi^{\tau,N}(\mathrm{d}v),\quad\phi\in C_{b}(H),$ (83)
@@
5.3 Strong law of large numbers
@@
Suppose that Assumptions 2-2 hold, then for any bounded uniformly continuous function $\phi:\ H\to\mathbb{R}$, the following strong law of large numbers holds
$\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\phi(u(t,\cdot))\mathrm{d}t=\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v),\quad\text{a.s.}$ (84)
Proof 5.12.
@@
$I_{p}(T)\leq\left(C\int_{0\leq s_{1}\leq s_{2}\leq T}\left(\mathbb{E}\left[\left|\hat{g}\left(s_{1},s_{2}\right)\right|^{p}\right]\right)^{\frac{1}{p}}\mathrm{~d}s_{1}\mathrm{d}s_{2}\right)^{p}.$ (85)
@@
$\begin{aligned} \left(\mathbb{E}\left[\left|\hat{g}\left(s_{1},s_{2}\right)\right|^{p}\right]\right)^{\frac{1}{p}}\leq&Ce^{-\frac{\epsilon_{0}}{2}(s_{2}-s_{1})}.\end{aligned}$ (86)
Combining (85) and (86), we obtain
@@
Taking the limit as $T\to\infty$, this completes the proof.
@@
Suppose that Assumptions 2, 2, 4.2 hold, then for any bounded uniformly continuous function $\phi:\ H\to\mathbb{R}$, the following strong law of large numbers holds
$\lim_{T\rightarrow\infty}\lim_{\tau\rightarrow 0}\lim_{N\rightarrow\infty}\frac{\tau}{T}\sum_{m=0}^{\lfloor T/\tau\rfloor}\phi(u^{\tau,N}({t_{m}},\cdot))=\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v),\quad\text{a.s.}$ (87)
Proof 5.14.
@@
$\frac{\tau}{T}\sum_{m=0}^{\lfloor T/\tau\rfloor}\phi(u^{\tau,N}({t_{m}},\cdot))-\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v)$ (88)
@@
This completes the proof.
6 Numerical example
@@
$\begin{cases}\dfrac{\partial u(t,x)}{\partial t}+\Delta^{2}u(t,x)=\frac{1}{2}\Delta\!\left(u^{3}(t,x)-u^{2}(t,x)+2u(t,x)-2\right)+\dfrac{\partial W(t,x)}{\partial t},&t\in(0,T],\ x\in(0,\pi),\\[6.0pt] \frac{\partial u(t,0)}{\partial x}=\frac{\partial u(t,\pi)}{\partial x}=\frac{\partial^{3}u(t,0)}{\partial x^{3}}=\frac{\partial^{3}u(t,\pi)}{\partial x^{3}}=0,&t\in(0,T],\\ u(0,x)=u_{0}(x),&x\in(0,\pi).\end{cases}$ (89)
@@
We first examine the strong convergence rates of the fully discrete scheme (54). The mean-square errors and strong convergence rates are reported in Table 1, which aligns with the theoretical predictions stated in Theorem 4.7. These results demonstrate that the mean-square temporal convergence rate of the full discretization is about $\frac{3}{8}$, and the spatial convergence rate is about $1$.
Table 1: Mean-square errors and strong convergence rates ( $T=100,$ $u_{0}=\frac{1}{3}\cos(x)+\frac{1}{3}$).
@@
Next, we investigate the long-time behavior of the solution to (89). According to Remark 5.10 and Theorem 5.13, the time average of $\phi\big(u^{\tau,N}(u_{0};t,\cdot)\big)$ converges almost surely to the ergodic limit $\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v)$, and t This motivates us to approximate the ergodic limit using two approaches:
@@
As shown in Figures 1 and 2, both Approximations I and II converge to the same ergodic limit, regardless of the initial condition. Table 2 compares the computational cost and error of both approximations. Notably, the CPU time required for Approximation II is significantly less than that for Approximation I. This efficiency improvement comes from the fact that Approximation II uses only a single sample trajectory, thereby avoiding the simulation of a large number of samples and offering an advantage in computational efficiency.
@@
Figure 1: The time average of $\mathbb{E}\left[\phi\big(u^{\tau,N}(u_{0};t,\cdot)\big)\right]$ started from different initial values.
@@
Figure 2: The time average of $\phi\big(u^{\tau,N}(u_{0};t,\cdot)\big)$ started from different initial values.
Table 2: Errors and CPU time for approximating the ergodic limit.
@@
Finally, we examine the dependence of the ergodic limit on initial conditions. We consider three distinct initial conditions, each belonging to a different $H_{\alpha}$ space: $u_{0}=\frac{1}{3}$ ($\alpha=\frac{1}{3}$), $\frac{1}{3}\cos(x)+1$ ($\alpha=1$), and $2\cos(2x)+\cos(x)+2$ ($\alpha=2$), together with the parameter triple $(v,\alpha_{1},\alpha_{2})=(e^{x},0,3)$. As shown in Figure 3, the ergodic limits will be different depending on the initial condition, which aligns with Remark 3.19 stating that the SCHE does not admit a unique invariant measure on $H$.
Figure 3: Ergodic limits for initial conditions in different spaces.

### v3

@@ 5.  Approximation of invariant measures
where $\nu(\cdot):=\lambda\left(\cdot\cap B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\cap D_{\alpha}\right)/\lambda\left(B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\cap D_{\alpha}\right).$ Letting $\eta=\gamma_{1}\gamma_{2}\lambda\left(B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\cap D_{\alpha}\right)$, we obtain $P_{k^{*}}(\vec{v},B\cap D_{\alpha})\geq\eta\nu(B\cap D_{\alpha})$ for all $B\in\mathscr{B}\big(\mathbb{R}^{N}\big)$ and $\vec{v}\in S\cap D_{\alpha}$. Since $B_{\epsilon_{2}}\left(\vec{w}^{*}\right)\subset S$, it follows that $\nu(S\cap D_{\alpha})=1$ and $\nu\left(\left(S\cap D_{\alpha}\right)^{c}\right)=0$. This completes the proof. ∎
Next we introduce the Lyapunov condition.
Assumption 5.3.
(Lyapunov condition) There is a function $V:\mathbb{R}^{N}\rightarrow[1,\infty)$ with $\lim_{|\vec{v}|\rightarrow\infty}V(\vec{v})=\infty$ and real numbers $\alpha_{1}\in$ $(0,1),\alpha_{2}\in[0,\infty)$ such that
@@
[31, Lemma A.6] Suppose $\{\vec{v}_{k}\}_{k\in\mathbb{N}}$ and $\{\vec{w}_{k}\}_{k\in\mathbb{N}}$ are two Markov chains satisfying Assumption 5.3. Let $\gamma\in(\sqrt{\alpha_{1}},1)$ and define
@@
By Lemmas 5.2 and 5.4, we can establish the following theorem.
@@
$\sup_{k\geq 0}\mathbb{E}\left[V\left(\vec{v}_{k}\right)\right]<\infty.$ (5.1)
@@
It follows from (5.1) that the sequence $\left\{\pi^{\tau,N}_{m}\right\}_{m\in\mathbb{N}}$ is tight. Consequently, by the Krylov–Bogoliubov Theorem [24, Theorem 1.2], the Markov chain $\left\{\vec{v}_{k}\right\}_{k\in\mathbb{N}}$ admits an invariant measure $\pi^{\tau,N}$ .
@@
Algorithm 1 Construction of the new Markov chain $\{\vec{w}_{k}\}$
@@
2:Output: Markov chain $\{\vec{w}_{k}\}$
@@
4:  Sample $\hat{\psi}_{k}$
5:  if $\vec{w}_{k}\in S$ then
6:   $\vec{w}_{k+1}\leftarrow\hat{\psi}_{k}\,\vec{w}_{k}^{*}+(1-\hat{\psi}_{k})\vec{\xi}_{k}$
7:  else
8:   $\vec{w}_{k+1}\leftarrow\vec{w}_{k}^{*}$
9:  end if
10:end for
@@
This shows that $\pi^{\tau,N}=\hat{\pi}^{\tau,N}$, and hence uniqueness is proved. ∎
@@
Suppose that Assumptions 2.2-2.4 hold, then $\left\{u^{\tau,N}(t_{m},\cdot)\right\}_{m\in\mathbb{N}}$ admits a unique invariant measure $\pi^{\tau,N}$.
Proof.
Since $u^{\tau,N}(t,x)=P_{N}U^{\tau,N}_{t}$, it is sufficient to establish the existence and uniqueness of an invariant measure for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$. From (4.9), it follows that $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ forms a time-homogeneous Markov chain. According to Theorem 5.1, establishing the existence and uniqueness of an invariant measure for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ necessitates verifying that this Markov chain satisfies both the Lyapunov and minorization conditions restricted to $D_{\alpha}$.
@@
for any $\vec{v}\in l^{2}_{N}$. By applying (4.9) and using properties of conditional expectation, we obtain
$\mathbb{E}\left[V\left(U^{\tau,N}_{t_{m+1}}\right)\mid\mathscr{F}_{t_{m}}\right]=$ $\sum_{j=1}^{N-1}\lambda_{N,j}^{-1}\mathbb{E}\left[\langle e^{-A_{N}^{2}\tau}U^{\tau,N}_{t_{m}}+\tau A_{N}e^{-A_{N}^{2}\tau}\tilde{F}_{N}(U^{\tau,N}_{t_{m}})+\sigma e^{-A_{N}^{2}\tau}\Delta\beta_{m},\phi_{N,j}\rangle^{2}_{l^{2}_{N}}\mid\mathscr{F}_{t_{m}}\right]$ (5.2)
@@
Substituting these inequalities into (5.2), we conclude that
@@
Now, for any $\vec{v}\in S\cap D_{\alpha}$, $\vec{v}^{*}\in\text{Int}(S)\cap D_{\alpha}$, $\delta>0$ and $\vec{w}\in B_{\delta}\left(\vec{v}^{*}\right)\cap D_{\alpha}$, we apply (4.9) together with the fact that $\left\{\phi_{N,j}\right\}_{j=0}^{N-1}$ is an orthonormal basis of $l^{2}_{N}$ to obtain
@@
This demonstrates that one can choose $\Delta\beta_{m}$ such that $U^{\tau,N}_{t_{m+1}}=\vec{w}$ starting from $U^{\tau,N}_{t_{m}}=\vec{v}$. Since $\left\{\left\langle\Delta\beta_{m},\phi_{N,j}\right\rangle_{l^{2}_{N}}\right\}_{j=1}^{N-1}$ are Gaussian random variables, it follows that $P_{1}\left(\vec{v},B_{\delta}\left(\vec{v}^{*}\right)\cap D_{\alpha}\right)>0$, which verifies the second part of the minorization condition. To verify the third condition in the minorization condition, observe that each Gaussian variable $\left\langle\Delta\beta_{m},\phi_{N,j}\right\rangle_{l^{2}_{N}}$ admits a $C^{\infty}$ density function for $j=1,\cdots,N-1$, and $\left\langle\Delta\beta_{m},\phi_{N,0}\right\rangle_{l^{2}_{N}}=0$. Hence, the one-step transition kernel $P_{1}(\vec{v},\cdot)$ admits a density $p_{1}(\vec{v},\vec{w})$, which is jointly continuous on $(\vec{v},\vec{w})\in\left(S\cap D_{\alpha}\right)\times\left(S\cap D_{\alpha}\right)$. Finally, the time-homogeneity of the Markov chain $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ implies that the density functions $\left\{p_{m}(\vec{v},\vec{w})\right\}_{m\in\mathbb{N}}$ are jointly continuous. This completes the proof. ∎
5.Approximation of invariant measure
@@
Suppose that Assumptions 2.2, 2.4, 4.5 hold, then for any bounded uniformly continuous function $\phi:\ H\to\mathbb{R}$, there exists a constant $C>0$, independent of $h$ and $\tau$, such that
$\left|\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v)-\int_{V^{N}}\phi(v)\tilde{\pi}^{\tau,N}(\mathrm{d}v)\right|\leq C\left(h^{1-\epsilon}+\tau^{\frac{3}{8}-\frac{\epsilon}{2}}\right),$ (5.3)
@@
$\lim_{M\rightarrow\infty}\frac{1}{M}\sum_{m=0}^{M-1}\mathbb{E}\left[\phi(u^{\tau,N}(t_{m},\cdot))\right]=\int_{V^{N}}\phi(v)\pi^{\tau,N}(\mathrm{d}v),\quad\phi\in C_{b}(H),$ (5.4)
@@
5.Strong law of large numbers
@@
Suppose that Assumptions 2.2-2.4 hold, then for any bounded uniformly continuous function $\phi:\ H\to\mathbb{R}$, the following strong law of large numbers holds
$\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\phi(u(t,\cdot))\mathrm{d}t=\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v),\quad\text{a.s.}$ (5.5)
Proof.
@@
$I_{p}(T)\leq\left(C\int_{0\leq s_{1}\leq s_{2}\leq T}\left(\mathbb{E}\left[\left|\hat{g}\left(s_{1},s_{2}\right)\right|^{p}\right]\right)^{\frac{1}{p}}\mathrm{~d}s_{1}\mathrm{d}s_{2}\right)^{p}.$ (5.6)
@@
$\begin{aligned} \left(\mathbb{E}\left[\left|\hat{g}\left(s_{1},s_{2}\right)\right|^{p}\right]\right)^{\frac{1}{p}}\leq&Ce^{-\frac{\epsilon_{0}}{2}(s_{2}-s_{1})}.\end{aligned}$ (5.7)
Combining (5.6) and (5.7), we obtain
@@
Taking the limit as $T\to\infty$, this completes the proof. ∎
@@
Suppose that Assumptions 2.2, 2.4, 4.5 hold, then for any bounded uniformly continuous function $\phi:\ H\to\mathbb{R}$, the following strong law of large numbers holds
$\lim_{T\rightarrow\infty}\lim_{\tau\rightarrow 0}\lim_{N\rightarrow\infty}\frac{\tau}{T}\sum_{m=0}^{\lfloor T/\tau\rfloor}\phi(u^{\tau,N}({t_{m}},\cdot))=\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v),\quad\text{a.s.}$ (5.8)
Proof.
@@
$\frac{\tau}{T}\sum_{m=0}^{\lfloor T/\tau\rfloor}\phi(u^{\tau,N}({t_{m}},\cdot))-\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v)$ (5.9)
@@
This completes the proof. ∎
6.  Numerical example
@@
$\begin{cases}\dfrac{\partial u(t,x)}{\partial t}+\Delta^{2}u(t,x)=\frac{1}{2}\Delta\!\left(u^{3}(t,x)-u^{2}(t,x)+2u(t,x)-2\right)+\dfrac{\partial W(t,x)}{\partial t},&t\in(0,T],\ x\in(0,\pi),\\[6.0pt] \frac{\partial u(t,0)}{\partial x}=\frac{\partial u(t,\pi)}{\partial x}=\frac{\partial^{3}u(t,0)}{\partial x^{3}}=\frac{\partial^{3}u(t,\pi)}{\partial x^{3}}=0,&t\in(0,T],\\ u(0,x)=u_{0}(x),&x\in(0,\pi).\end{cases}$ (6.1)
@@
We first examine the strong convergence rates of the fully discrete scheme (4.8). The mean-square errors and strong convergence rates are reported in Table 6.1, which aligns with the theoretical predictions stated in Theorem 4.2. These results demonstrate that the mean-square temporal convergence rate of the full discretization is about $\frac{3}{8}$, and the spatial convergence rate is about $1$.
Table 6.1: Mean-square errors and strong convergence rates ( $T=100,$ $u_{0}=\frac{1}{3}\cos(x)+\frac{1}{3}$).
@@
Next, we investigate the long-time behavior of the solution to (6.1). According to Remark 5.5 and Theorem 5.5, the time average of $\phi\big(u^{\tau,N}(u_{0};t,\cdot)\big)$ converges almost surely to the ergodic limit $\int_{H_{\alpha}}\phi(v)\pi(\mathrm{d}v)$, and t This motivates us to approximate the ergodic limit using two approaches:
@@
As shown in Figures 6.1 and 6.2, both Approximations I and II converge to the same ergodic limit, regardless of the initial condition. Table 6.2 compares the computational cost and error of both approximations. Notably, the CPU time required for Approximation II is significantly less than that for Approximation I. This efficiency improvement comes from the fact that Approximation II uses only a single sample trajectory, thereby avoiding the simulation of a large number of samples and offering an advantage in computational efficiency.
@@
Figure 6.1: The time average of $\mathbb{E}\left[\phi\big(u^{\tau,N}(u_{0};t,\cdot)\big)\right]$ started from different initial values.
@@
Figure 6.2: The time average of $\phi\big(u^{\tau,N}(u_{0};t,\cdot)\big)$ started from different initial values.
Table 6.2: Errors and CPU time for approximating the ergodic limit.
@@
Finally, we examine the dependence of the ergodic limit on initial conditions. We consider three distinct initial conditions, each belonging to a different $H_{\alpha}$ space: $u_{0}=\frac{1}{3}$ ($\alpha=\frac{1}{3}$), $\frac{1}{3}\cos(x)+1$ ($\alpha=1$), and $2\cos(2x)+\cos(x)+2$ ($\alpha=2$), together with the parameter triple $(v,\alpha_{1},\alpha_{2})=(e^{x},0,3)$. As shown in Figure 6.3, the ergodic limits will be different depending on the initial condition, which aligns with Remark 3.8 stating that the SCHE does not admit a unique invariant measure on $H$.
Figure 6.3: Ergodic limits for initial conditions in different spaces.